The ratio of linear expansivity of copper to that of iron Is approximately 1:5. A specimen of iron and specimen of copper expand by the same amount per unit rise in temperature. What is the ratio of the length of iron to that of copper ?


1
Expert's answer
2021-05-27T17:47:48-0400

The expansion per unit rise in temperature is given as follows:


"\dfrac{\Delta L}{\Delta T} = \alpha L"

where "\Delta L" is the increase in length, "\Delta T" is the rise in temperature, "\alpha" is the expansivity, and "L" is the initial length.

According to the text of the problem, "\left( \dfrac{\Delta L}{\Delta T} \right)_{cooper } = \left( \dfrac{\Delta L}{\Delta T} \right)_{iron }" and "\dfrac{\alpha_{cooper}}{\alpha_{iron}} = \dfrac15". Then the ration of the lengths is:


"L_{cooper} = \dfrac{1}{\alpha_{cooper}} \left( \dfrac{\Delta L}{\Delta T} \right)_{cooper } \\
L_{iron} = \dfrac{1}{\alpha_{iron}} \left( \dfrac{\Delta L}{\Delta T} \right)_{iron }\\
\dfrac{L_{iron} }{L_{cooper} } = \dfrac{\dfrac{1}{\alpha_{iron}} \left( \dfrac{\Delta L}{\Delta T} \right)_{iron }}{ \dfrac{1}{\alpha_{cooper}} \left( \dfrac{\Delta L}{\Delta T} \right)_{cooper }} = \dfrac{\alpha_{cooper}}{\alpha_{iron}} = \dfrac15"

Answer. 1:5.